OpenStudy (anonymous):

In a certain lottery game, you must pick five digits in order from 0 to 9 and duplicates are allowed. If you win, the prize is $10,000. What lottery ticket price would make this game fair?

OpenStudy (anonymous):

The number of ways to pick digits is: \[ 10^5 \]The number of ways to win is \(1\). It's fair if the expected value is \(0\). \[ E[X] = 0 = 10,000\frac{1}{10^5} + x\frac{10^5-1}{10^5} \]

OpenStudy (anonymous):

Whoops, that formula needs to be changed....

OpenStudy (anonymous):

\[ E[X] = 0 = (10,000-x)\frac{1}{10^5} + (-x)\frac{10^5-1}{10^5} \]

OpenStudy (anonymous):

Now you'd solve for \(x\).

OpenStudy (anonymous):

A $0.17 B $0.24 C $0.08 D $0.10 ???

OpenStudy (anonymous):

.10??

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

Thanks ! :)